thinfilms morning report 2/16/05

james haydon jhaydon at snf.stanford.edu
Wed Feb 16 11:24:50 PST 2005


Thinfilms Morning Report 2/16/2005

 

 

Gryphon:            PPC not working hope to have it repaired by early 
next week

Innotec;            ok

STS Dep:            ok

Implanter:            having vacuum and high voltage problems should be 
able to get to it late next week

Metalica:            ok

SCT:                ok

CPD:                ok                   

WbGen-CTB:            ok

WbGen-RHP:            ok

WbGen:-LHP:            ok

WbGaAs-RHP:            ok

WbGaSa-LHP:            ok

WbSilicide:            ok

WbDif: ok

WbNonmetal:            ok

WbNitride:            ok

WbMetal:            ok

Wbmiscres:            ok

WBOptres:            ok

WBSolvent:            ok

DI water:            ok

LN:                  ok

 

-------------- next part --------------
An HTML attachment was scrubbed...
URL: <http://snf.stanford.edu/pipermail/thinfilms/attachments/20050216/0a85bd18/attachment.html>


More information about the thinfilms mailing list